When purchasing bulk orders of​ batteries, a toy manufacturer uses this acceptance sampling​ plan: Randomly select and test 41 batteries and determine whether each is within specifications. The entire shipment is accepted if at most 3 batteries do not meet specifications. A shipment contains 4000 ​batteries, and 2​% of them do not meet specifications. What is the probability that this whole shipment will be​ accepted? Will almost all such shipments be​ accepted, or will many be​ rejected?


1
Expert's answer
2021-03-15T08:42:28-0400

Let "X=" the number of defective batteries. The probability distribution of the random variable "X," hypergeometric distribution, is given by

"h(x;n,M,N)= \dfrac{\dbinom {M}{x}\dbinom{N-M}{n-x}}{\dbinom{N}{n}}"

Given that "N=4000, M=4000\cdot0.02=80,n=41."

What is the probability that at most 3 batteries do not meet specifications?


"P(X\leq3)=P(X=0)+P(X=1)+""+P(X=2)+P(X=3)="




"={\binom{80}{0}\binom{4000-80}{41-0} \over \binom{4000}{41}} +{\binom{80}{1}\binom{4000-80}{41-1} \over \binom{4000}{41}}+""+{\binom{80}{2}\binom{4000-80}{41-2} \over \binom{4000}{41}}+{\binom{80}{3}\binom{4000-80}{41-3} \over \binom{4000}{41}}\approx"


"\approx 0.4349+0.3676+0.1497+0.0390=0.9913"

The probability that this whole shipment will be accepted is  "0.9913."